OpenStudy (anonymous):

Solve by factoring x^2+17x+72=0

OpenStudy (anonymous):

not poss

OpenStudy (anonymous):

what multiplies to be 72 and adds to 17? I don't know

OpenStudy (anonymous):

8 x 9 = 72 8 + 9 = 17

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(x+8)(x+9) = 0 (x+8) = 0 (x+9) = 0 x = -8, x = -9

OpenStudy (anonymous):

yes